OpenStudy (anonymous):

HELP! Find the product of z1=5(cos 20 degrees+ i sin 20 degrees) and z2=4(cos 10 degrees + i sin 10 degrees) and write the answer in standard form.

OpenStudy (anonymous):

i really need help fast

jimthompson5910 (jim_thompson5910):

Hint: If z1 = r1*cis(theta1) and z2 = r2*cis(theta2) Then z1*z2 = (r1*r2)*cis(theta1+theta2)

OpenStudy (apoorvk):

great, just convert them into euler's form, that is: \(re^{i\theta} = r(\cos \theta + \iota \sin \theta)\) And then multiply both the complex nos. this should be the easiest method.

OpenStudy (anonymous):

can you show me?please

jimthompson5910 (jim_thompson5910):

In z1=5(cos 20 degrees+ i sin 20 degrees) , what is r and theta?

OpenStudy (apoorvk):

well, your two nos. are: \(z_1 = 5(\cos 20^0 + \iota \sin 20^0) = 5 e^{i 20^o}\) Similarly what would \(z_2\) be?

OpenStudy (anonymous):

is it 6.4

OpenStudy (anonymous):

Can u help

jimthompson5910 (jim_thompson5910):

Not sure how you're getting 6.4

OpenStudy (anonymous):

I solved that equation? can u walk through please? i did it wrong

jimthompson5910 (jim_thompson5910):

In z1=5(cos 20 degrees+ i sin 20 degrees), r1 is 5 and theta1 is 20 In z2=4(cos 10 degrees + i sin 10 degrees), r2 is 4 and theta2 is 10 So z1*z2 = (r1*r2)*cis(theta1+theta2) z1*z2 = (5*4)*cis(20+10) z1*z2 = 20*cis(30) z1*z2 = 20*( cos(30) + i*sin(30) ) z1*z2 = 20*( sqrt(3)/2 + i*(1/2) ) z1*z2 = 20*( sqrt(3)/2) +20*( i*(1/2) ) z1*z2 = 10*sqrt(3) + 10i

jimthompson5910 (jim_thompson5910):

The first part is just recognizing/remembering the formula. The remaining parts involve simplifying really.